# Strong order-one convergence of the Euler method for random ordinary differential equations driven by semi-martingale noises

### v7

@@ 8.1. Estimating the order of convergence
For estimating the order of convergence, we use the Monte Carlo method, computing a number of numerical approximations $\{X_{t_{j}}^{N_{i}}(\omega_{m})\}_{j=0,\ldots,N_{i}}$, of sample path solutions $\{X_{t}(\omega_{m})\}_{t\in I}$, for samples $\omega_{m},$ with $m=1,\ldots,M$, and mesh resolutions $N_{i},$ with $i=1,\ldots,I,$ and taking the maximum in time of the average of their absolute differences at the mesh points:
@@
Then we fit the errors $\epsilon^{N_{i}}$ to the power law $C\Delta t_{N_{i}}^{p}$, in order to find $p$, along with a 95% confidence interval.
@@
    A series of time steps $\Delta t_{N_{i}}=T/N_{i}$, with $N_{i}=2^{n_{i}},$ for some $n_{i}\in\mathbb{N},$ $i=1,\ldots,I,$ so that finer meshes are refinements of coarser meshes.
@@
    We fit $(\Delta t_{N_{i}},\epsilon^{N_{i}})$ to the power law $C\Delta t_{N_{i}}^{p}$, via linear least-square regression in log scale, so that $\ln\epsilon^{N_{i}}\approx\ln C+p\ln\Delta t_{N_{i}}$, for suitable $C$ and $p$, with $p$ giving the order of convergence. This amounts to solving the normal equation $(A^{t}A)\mathbf{v}=A^{t}\ln(\boldsymbol{\epsilon})$, where $\mathbf{v}$ is the vector $\mathbf{v}=(\ln(C),p)$, $A$ is the Vandermonde matrix associated with the logarithm of the mesh steps $(\Delta t_{N_{i}})_{i}$, and $\ln(\boldsymbol{\epsilon})$ is the vector $\ln(\boldsymbol{\epsilon})=(\ln(\epsilon^{N_{i}}))_{i}$.
@@
    and compute the 95% confidence interval $[\epsilon_{\min},\epsilon_{\max}]$ for the strong error with
    $\epsilon_{\min}^{N_{i}}=\max_{j=0,\ldots,N_{i}}(\epsilon_{t_{j}}^{N_{i}}-2\sigma_{t_{j}}^{N_{i}}),\quad\epsilon_{\max}^{N_{i}}=\max_{j=0,\ldots,N_{i}}(\epsilon_{t_{j}}^{N_{i}}+2\sigma_{t_{j}}^{N_{i}}).$
@@
    Finally, from the normal equations above, we compute the 95% confidence interval $[p_{\min},p_{\max}]$ by computing the minimum and maximum values of $p$ in the image, by the linear map $\mathbf{e}\mapsto(\ln(C),p)=(A^{t}A)^{-1}A^{t}\ln(\mathbf{e})$, of the polytope formed by all combinations of the indices in $(\epsilon_{\min}^{N_{i}})_{i}$ and $(\epsilon_{\max}^{N_{i}})_{i}$, which is the image of the set of 95% confidence intervals for the errors obtained with the Monte-Carlo approximation of the strong error.
8.2. Numerical setup
@@ 8.2.1. Simulating a Random ODE
$\begin{cases}\displaystyle\frac{\mathrm{d}X_{t}}{\mathrm{d}t}=f(t,X_{t},Y_{t}),&0\leq t\leq T,\\ \left.X_{t}\right|_{t=0}=X_{0},\end{cases}$ (8.2)
@@
8.2.2. Calculating the order of convergence
@@
    suite = ConvergenceSuite(t0, tf, x0law, f, noise,params, target, method, ntgt, ns, m, $\texttt{suite = ConvergenceSuite(t0, tf, x0law, f, noise,}\\ \texttt{params, target, \allowbreak method, \allowbreak ntgt, \allowbreak ns, \allowbreak m, ks)}.$
@@
8.3. Numerical examples
We now describe in more details the examples used to illustrate the order of convergence. We start with some linear equations with all sorts of noises. Then we illustrate the $\min\{H+1/2,1\}$ order of convergence in the case of a fractional Brownian motion (fBm) noise with Hurst parameter $10<H<$ 1 . Next we move to a number of linear and nonlinear models, as mentioned in the Section 1.
8.3.1. Homogeneous linear equation with Wiener noise
@@
$\begin{cases}\displaystyle\frac{\mathrm{d}X_{t}}{\mathrm{d}t}=W_{t}X_{t},\qquad 0\leq t\leq T,\\ \left.X_{t}\right|_{t=0}=X_{0}.\end{cases}$ (8.3)
@@
Equation Equation 8.3 has the explicit solution
$X_{t}=e^{\int_{0}^{t}W_{s}\;\mathrm{d}s}X_{0}.$ (8.4)
@@
The exact distribution of $\int_{0}^{\tau}W_{s}\;\mathrm{d}s$ given the step $\Delta W=W_{\tau}-W_{0}=W_{\tau}$ is computed in [18, Section 14.2] as
$\int_{0}^{\tau}W_{s}\;\mathrm{d}s\sim\frac{\tau}{2}\Delta W+\sqrt{\frac{\tau^{3}}{12}}\mathcal{N}(0,1)=\frac{\tau}{2}\Delta W+\mathcal{N}\left(0,\frac{\tau^{3}}{12}\right).$ (8.5)
As for the distribution of the integral over a mesh interval $[t_{i},t_{i+1}]$ when given the endpoints $W_{t_{i}}$ and $W_{t_{i+1}},$ we use that $s\mapsto W_{t_{i}+s}-W_{t_{i}}$ is a standard Wiener process to find, from Equation 8.5, that
@@
$Z_{i}\sim\mathcal{N}\left(0,\frac{(t_{i+1}-t_{i})^{3}}{12}\right)=\frac{\sqrt{(t_{i+1}-t_{i})^{3}}}{\sqrt{12}}\mathcal{N}(0,1).$ (8.6)
@@
$\int_{0}^{t_{j}}W_{s}\;\mathrm{d}s=\sum_{i=0}^{j-1}\int_{t_{i}}^{t_{i+1}}W_{s}\;\mathrm{d}s=\sum_{i=0}^{j-1}\left(\frac{(W_{t_{i+1}}+W_{t_{i}})}{2}(t_{i+1}-t_{i})+Z_{i}\right),$ (8.7)
where $Z_{0},\ldots,Z_{N-1}$ are independent random variables with distribution Equation 8.6.
Notice the first term in Equation 8.7 is the trapezoidal rule, with the second one being a correction term.
@@
Table 1. Mesh points (N), time steps (dt), strong error (error), and standard error (std err) of the Euler method for $\mathrm{d}X_{t}/\mathrm{d}t=W_{t}X_{t}$ for each mesh resolution $N$, with initial condition $X_{0}\sim\mathcal{N}(0,1)$ and a standard Wiener process noise $\{W_{t}\}_{t}$, on the time interval $I=[0.0,1.0]$, based on $M=200$ sample paths for each fixed time step, with the target solution calculated with $65536$ points. The order of strong convergence is estimated to be $p=1.001$, with the 95% confidence interval $[0.936,1.0664]$.
Hence, once an Euler approximation of Equation 8.3 is computed, along with realizations $\{W_{t_{i}}\}_{i=0}^{N}$ of a sample path of the noise, we consider an exact pathwise solution given by
$X_{t_{j}}=X_{0}e^{\sum_{i=0}^{j-1}\left(\frac{1}{2}\left(W_{t_{i}}+W_{t_{i+1}}\right)(t_{i+1}-t_{i})+Z_{i}\right)},$ (8.9)
for realizations $Z_{i}$ drawn from a normal distributions given by Equation 8.6. Figure 7 shows an approximate solution and a few sample paths of exact solutions associated with the given realizations of the noise on the mesh points.
The function $f(t,x,y)=yx$ associated with equation Equation 8.3 does not satisfy the uniform Lipschitz condition Equation 2.6. Nevertheless, thanks to the form of the exact solution Equation 8.4 the form of the approximation Equation 8.9 and estimates such as Equation 8.8, the proof of Theorem 7.1 can be adapted to show strong order $1$ convergence in this case, as well.
@@
8.3.2. Non-homogeneous linear system of RODEs with different types of noises
Next, we examine a linear equation with noise present in both the homogeneous and nonhomogeneous terms, approached in two ways: first, as a series of scalar equations, each with a unique univariate noise in a family of different types of noises; and second, as a system of equations that combines all these noise types. For most of the noises, the current knowledge expects a lower order of strong convergence than the strong order 1 we prove here. The aim of this section is to illustrate this improvement for all such noises.
The equations take the form
where $\{{X}_{t}\}_{t}$ is either a scalar or a vector-valued process, and $\{{Y}_{t}\}_{t}$ is, correspondingly, either a scalar or a vector-valued noise process with the same dimension as ${X}_{t}$. In the case of the system, each coordinate of $\{{Y}_{t}\}_{t}$ is a scalar noise process independent of the noises in the other coordinates. The scalar noises used in the simulations are the following, in the order of coordinates of ${Y}_{t}$:
  1. (i)
    A standard Wiener process (W);
  2. (ii)
  3. (iii)
  4. (iv)
    A non-autonomous homogeneous linear Itô process (hlp) $\{H_{t}\}_{t}$ given by $\mathrm{d}H_{t}=(\mu_{1}+\mu_{2}\sin(\vartheta t))H_{t}\;\mathrm{d}t+\sigma\sin(\vartheta t)H_{t}\;\mathrm{d}W_{t}$ with $\mu_{1}=0.5,$ $\mu_{2}=0.3,$ $\sigma=0.5,$ $\vartheta=3\pi,$ and initial condition $H_{0}=0.2;$
  5. (v)
  6. (vi)
  7. (vii)
  8. (viii)
  9. (ix)
N dt error std err
64 0.0156 0.209 0.0198
128 0.00781 0.102 0.00973
256 0.00391 0.0498 0.00473
512 0.00195 0.0251 0.00239
Table 2. Mesh points (N), time steps (dt), strong error (error), and standard error (std err) of the Euler method for $\mathrm{d}{X}_{t}/\mathrm{d}t=-\|{Y}_{t}\|^{2}{X}_{t}+{Y}_{t}$ for each mesh resolution $N$, with initial condition ${X}_{0}\sim\mathcal{N}({0},\mathrm{I})$ and vector-valued noise $\{{Y}_{t}\}_{t}$ with all the implemented noises, on the time interval $I=[0.0,1.0]$, based on $M=92$ sample paths for each fixed time step, with the target solution calculated with $262144$ points. The order of strong convergence is estimated to be $p=1.021$, with the 95% confidence interval $[0.9108,1.1308]$.
Figure 9. Order of convergence of the strong error of the Euler method for $\mathrm{d}{X}_{t}/\mathrm{d}t=-\left\|{Y}_{t}\right\|^{2}{X}_{t}+{Y}_{t},$ based on Table 2.
Figure 11. Order of convergence for the equation Equation 8.10 as it varies among different types of noises, starting with all the noises combined in a system of equations (all) and continuing with the scalar cases, each with a different type of noise, according to the table Table 3
noise $p$ $p_{\textrm{min}}$ $p_{\textrm{max}}$
all noises combined (all) 1.02082 0.910835 1.13082
Wiener (W) 1.00851 0.893154 1.11785
OrnsteinUhlenbeck (OU) 0.957151 0.854775 1.06022
GeometricBrownianMotion (gBm) 0.925382 0.770407 1.08201
HomogeneousLinearIto (hlp) 1.00657 0.878322 1.13411
CompoundPoisson (cP) 0.975898 0.735864 1.22554
PoissonStep (sP) 0.969316 0.851299 1.08621
ExponentialHawkes (H) 1.0213 0.92944 1.11311
Transport (T) 1.06509 0.983371 1.14687
FractionalBrownianMotion (fBm) 1.01897 0.887446 1.14729
Table 3. Type of noise, order $p$ of strong convergence, and 95% confidence interval $p_{\textrm{min}}$ and $p_{\textrm{max}}$, for the linear equation Equation 8.10 with various types of noise.
The strong order 1 convergence is not a surprise in the case of the Wiener and Ornstein-Uhlenbeck process since the corresponding RODE can be turned into an SDE with an additive noise. In this case, the Euler-Maruyama approximation for the noise part of the SDE is distributionally exact, and the Euler method for the RODE becomes equivalent to the Euler-Maruyama method for the SDE, and it is known that the Euler-Maruyama method for an SDE with additive noise is of strong order 1 [19]. The geometric Brownian motion is also expected to yield strong order 1 convergence thanks to the results in [39] and the fact that it is an Itô diffusion process. For the remaining noises, however, previous works would estimate the order of convergence to be below the order 1 attained here.
Notice we chose the Hurst parameter of the fractional Brownian motion process to be between 1/2 and 1, so that our results guarantee that the strong convergence is also of order 1, just like for the other types of noises in $\{{Y}_{t}\}_{t}$. Previous knowledge would expect a strong convergence of order $H$, with $1/2<H<1,$ instead.
8.3.3. Fractional Brownian motion noise

### v8

@@ 8.1. Estimating the order of convergence
For estimating the order of convergence, we use the Monte Carlo method, computing a number of numerical approximations $\{X_{t_{j}}^{N_{i}}(\omega_{m})\}_{j=0,\ldots,N_{i}}$, of sample path solutions $\{X_{t}(\omega_{m})\}_{t\in I}$, for samples $\omega_{m},$ with $m=1,\ldots,M$, and mesh resolutions $N_{i},$ with $i=1,\ldots,i_{\max},$ and taking the maximum in time of the average of their absolute differences at the mesh points:
@@
Then we fit the errors $\epsilon^{N_{i}}$ to the power law $C\Delta t_{N_{i}}^{p}$, in order to find $p$, along with a 95% confidence interval. The fitting process and the estimation of the confidence intervals follow classical methods in Statistics [14, 21, 25, 27]. Reporting confidence intervals in the context of approximations of stochastic differential equations, in particular, can also be seen in [32].
@@
    A series of time steps $\Delta t_{N_{i}}=T/N_{i}$, with $N_{i}=2^{n_{i}},$ for some $n_{i}\in\mathbb{N},$ $i=1,\ldots,i_{\max},$ so that finer meshes are refinements of coarser meshes.
@@
    We fit $(\Delta t_{N_{i}},\epsilon^{N_{i}})$ to the power law $C\Delta t_{N_{i}}^{p}$, via linear least-square regression in log scale, so that $\ln\epsilon^{N_{i}}\approx\ln C+p\ln\Delta t_{N_{i}}$, for suitable $C$ and $p$, with $p$ giving the order of convergence. This amounts to solving the normal equation $(A^{\operatorname{tr}}A)\mathbf{v}=A^{\operatorname{tr}}\ln(\boldsymbol{\epsilon})$, where $\mathbf{v}$ is the vector $\mathbf{v}=(\ln(C),p)$, $A$ is the Vandermonde matrix associated with the logarithm of the mesh steps $(\Delta t_{N_{i}})_{i}$, and $\ln(\boldsymbol{\epsilon})$ is the vector $\ln(\boldsymbol{\epsilon})=(\ln(\epsilon^{N_{i}}))_{i}$.
@@
    and compute the 95% confidence interval $\epsilon_{\min}^{{N_{i}}},\epsilon_{\max}^{{N_{i}}}]$ for the strong error at each mesh resolution $N_{i}$ with
    $\epsilon_{\min}^{N_{i}}=\max_{j=0,\ldots,N_{i}}(\epsilon_{t_{j}}^{N_{i}}-{1.96}\sigma_{t_{j}}^{N_{i}}),\quad\epsilon_{\max}^{N_{i}}=\max_{j=0,\ldots,N_{i}}(\epsilon_{t_{j}}^{N_{i}}+{1.96}\sigma_{t_{j}}^{N_{i}}).$
@@
    Next we estimate the 95% confidence interval for the joint distribution $\boldsymbol{\epsilon}=(\epsilon^{N_{i}})_{i=1,\ldots,i_{\max}}$ using the Bonferroni inequality [27]. More precisely, we find the $z$-score value for the $\alpha=1-0.05/i_{\max}$ confidence level for a univariate normal distribution and use that to obtain the $\alpha$ confidence intervals $\epsilon_{\alpha,\min}^{N_{i}},\epsilon_{\alpha,\max}^{N_{i}}]$ for the marginals $\epsilon^{N_{i}}$ with
    $\epsilon_{\alpha,\min}^{N_{i}}=\max_{j=0,\ldots,N_{i}}(\epsilon_{t_{j}}^{N_{i}}-z\sigma_{t_{j}}^{N_{i}}),\quad\epsilon_{\alpha,\max}^{N_{i}}=\max_{j=0,\ldots,N_{i}}(\epsilon_{t_{j}}^{N_{i}}+z\sigma_{t_{j}}^{N_{i}}).$
    Then, the probability of the region $\Pi_{i=1}^{i_{\max}}[\epsilon_{\alpha,\min}^{N_{i}},\epsilon_{\alpha,\max}^{N_{i}}]$ is bounded from below by $1-i_{\max}(1-\alpha)=1-0.05=0.95,$ which means this is the confidence region of confidence level (at least) 95%, for $\boldsymbol{\epsilon}$.
  9. (ix)
    Finally, from the normal equations and the confidence region for $\boldsymbol{\epsilon}$, we compute the (at least) 95% confidence interval $p_{\min},p_{\max}]$ for $p$ via the transformation method. We do so by computing the minimum and maximum values of $p$ in the image of the confidence region $\Pi_{i=1}^{i_{\max}}[\epsilon_{\alpha,\min}^{N_{i}},\epsilon_{\alpha,\max}^{N_{i}}]$ by the map $\mathbf{e}\mapsto(\ln(C),p)=(A^{\operatorname{tr}}A)^{-1}A^{\operatorname{tr}}\ln(\mathbf{e})$. The map $\mathbf{e}\mapsto\ln(\mathbf{e})$ takes the hyperrectangle $\Pi_{i=1}^{i_{\max}}[\epsilon_{\alpha,\min}^{N_{i}},\epsilon_{\alpha,\max}^{N_{i}}]$ into the hyperrectangle $\Pi_{i=1}^{i_{\max}}[\ln(\epsilon_{\alpha,\min}^{N_{i}}),\ln(\epsilon_{\alpha,\max}^{N_{i}})].$ This hyperrectangle is taken onto a polygon in the plane $(C,p)$ by the linear map $(A^{\operatorname{tr}}A)^{-1}A^{\operatorname{tr}}$}, which is then projected onto an interval $p_{\min},p_{\max}]$ in the $p$ axis. Since the pseudo-inverse $(A^{\operatorname{tr}}A)^{-1}A^{\operatorname{tr}}$} and the projection are not bijective, the corresponding confidence level of this interval is at least 95%, but probably higher. In order to find $p_{\min}$ and $p_{\max},$ we just need to look at the vertices of the polygon in $(C,p)$, which are the set of points obtained by all combinations of the indices in $(\epsilon_{{\alpha,}\min}^{N_{i}})_{i}$ and $(\epsilon_{{\alpha,}\max}^{N_{i}})_{i}.$ We take the minimum of the $p$-coordinate of such points to be $p_{\min}$ and the maximum to be $p_{\max}.$
8.2. Non-homogeneous linear system of RODEs with different types of noises
We examine a linear equation with noise present in both the homogeneous and nonhomogeneous terms, approached in two ways: first, as a series of scalar equations, each with a unique univariate noise in a family of different types of noises; and secondly, as a system of equations that combines all these noise types. For most of the noises, the current knowledge expects a lower order of strong convergence than the strong order 1 we prove here. The aim of this section is to illustrate this improvement for all such noises.
The equations take the form
where $\{X_{t}\}_{t\in I}$ is either a scalar or a vector-valued process, and $\{Y_{t}\}_{t\in I}$ is, correspondingly, either a scalar or a vector-valued noise process with the same dimension as ${X}_{t}$. In the case of a system, each coordinate of $\{Y_{t}\}_{t\in I}$ is a scalar noise process independent of the noises in the other coordinates. The scalar noises used in the simulations are the following, in the order of coordinates of $Y_{t}$:
  1. (i)
    A standard Wiener process (W);
  2. (ii)
  3. (iii)
  4. (iv)
    A non-autonomous homogeneous linear Itô process (hlp) $\{H_{t}\}_{t\in I}$ given by $\mathrm{d}H_{t}=(\mu_{1}+\mu_{2}\sin(\vartheta t))H_{t}\;\mathrm{d}t+\sigma\sin(\vartheta t)H_{t}\;\mathrm{d}W_{t}$ with $\mu_{1}=0.5,$ $\mu_{2}=0.3,$ $\sigma=0.5,$ $\vartheta=3\pi,$ and initial condition $H_{0}=0.2;$
  5. (v)
  6. (vi)
  7. (vii)
  8. (viii)
  9. (ix)
N dt error std err
64 0.0156 0.205 0.0093
128 0.00781 0.0993 0.00451
256 0.00391 0.049 0.00223
512 0.00195 0.0243 0.00111
Table 1. Mesh points (N), time steps (dt), strong error (error), and standard error (std err) of the Euler method for $\mathrm{d}{X}_{t}/\mathrm{d}t=-\|{Y}_{t}\|^{2}{X}_{t}+{Y}_{t}$ for each mesh resolution NN, with initial condition ${X}_{0}\sim\mathcal{N}({0},\mathrm{I})$ and vector-valued noise $\{{Y}_{t}\}_{t}$ with all the implemented noises, on the time interval $I=[0.0,1.0]$, based on $M=400$ sample paths for each fixed time step, with the target solution calculated with $262144$ points. The order of strong convergence is estimated to be $p=1.024$, with the 95% confidence interval $[0.8926,1.1556]$.
Figure 7. Estimated order of convergence of the strong error of the Euler method for $\mathrm{d}{X}_{t}/\mathrm{d}t=-\left\|{Y}_{t}\right\|^{2}{X}_{t}+{Y}_{t},$ based on Table 1.
Figure 9. Estimated order of convergence for the equation Equation 8.2 as it varies among different types of noises, starting with all the noises combined in a system of equations (all) and continuing with the scalar cases, each with a different type of noise, according to Table 2.
noise $p$ $p_{\textrm{min}}$ $p_{\textrm{max}}$
all noises combined (all) 1.02409 0.892626 1.15557
Wiener (W) 1.0005 0.860131 1.1361
OrnsteinUhlenbeck (OU) 0.98683 0.858187 1.11547
GeometricBrownianMotion (gBm) 0.990724 0.806347 1.17738
HomogeneousLinearIto (hlp) 0.989188 0.825703 1.15296
CompoundPoisson (cP) 1.01923 0.742455 1.2723
PoissonStep (sP) 0.98615 0.835225 1.1373
ExponentialHawkes (H) 1.02111 0.910741 1.13141
Transport (T) 1.0677 0.97865 1.15692
FractionalBrownianMotion (fBm) 0.983487 0.840285 1.12595
Table 2. Type of noise, order $p$ of strong convergence, and 95% confidence interval endpoints $p_{\textrm{min}}$ and $p_{\textrm{max}}$, for the linear equation Equation 8.2 with various types of noise.
The strong order 1 convergence is not a surprise in the case of the Wiener and Ornstein-Uhlenbeck process since the corresponding RODE can be turned into an SDE with an additive noise. In such a case, the Euler-Maruyama approximation for the noise part of the SDE is distributionally exact, and the Euler method for the RODE becomes equivalent to the Euler-Maruyama method for the SDE, and it is known that the Euler-Maruyama method for an SDE with additive noise is of strong order 1 [23]. The geometric Brownian motion is also expected to yield strong order 1 convergence thanks to the results in [48] and the fact that it is an Itô diffusion process. For the remaining noises, however, previous works would show the order of convergence to be below the order 1 attained here.
Notice we chose the Hurst parameter of the fractional Brownian motion process to be between 1/2 and 1, so that our results for fBm indicate that the strong convergence should also be of order 1, just like for the other types of noises in $\{Y_{t}\}_{t\in I}$. Previous knowledge would expect a strong convergence of order $H$, with $1/2<H<1,$ instead.
Data Availability Statement
The research code associated with this article is available at the open-access, github repository https://github.com/rmsrosa/rode_convergence_euler, and detailed documentation for the repository is available at the website https://rmsrosa.github.io/rode_convergence_euler [30].
Acknowledgments
The authors thank Glauco Valle (IM/UFRJ), for helpful discussions about semi-martingales, Luan Lima Freitas (IM/UFRJ), for many helpful comments on the manuscript, and Ralph dos Santos Silva (IM/UFRJ) for helpful discussions about the estimates for the confidence intervals. We also like to thank the reviewers for a careful reading of the manuscript, with several comments and suggestions that helped improve the paper. The second author was partly supported by Laboratório de Matemática Aplicada, Instituto de Matemática, Universidade Federal do Rio de Janeiro (LabMA/ IM/UFRJ), Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES), Brasil, grant 001, and Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Brasília, Brasil, grant no. 408751/2023-1.
Appendix A Numerical setup and examples
In this supplementary material, we describe in details the remaining numerical examples illustrating the main work. In all the examples driven by semi-martingale noises, we numerically verify the strong order 1 convergence of the Euler method. In the example with fractional Brownian motion, we verify the strong convergence of order $\min\{H+1/2,1\},$ where $H$ is the Hurst parameter. The order of convergence is displayed with a 95% confidence interval that perfectly matches the theory developed in the paper. These estimates show that the theoretical estimate is sharp.
One of the distinguishing powers of the Julia language is to work both as a high-level, interactive, scripting language (with its dynamically typed and just-in-time compilation system), and as a fast, high-performance language (with its aggressive compilation capabilities). It allows for both easy prototyping and the production of efficient end code. As such, it is growing quickly as a language suitable for scientific computing and computationally-demanding real-world applications.
A.1. Numerical setup
@@ A.1.1. Simulating a Random ODE
$\begin{cases}\displaystyle\frac{\mathrm{d}X_{t}}{\mathrm{d}t}=f(t,X_{t},Y_{t}),&0\leq t\leq T,\\ \left.X_{t}\right|_{t=0}=X_{0},\end{cases}$ (A.1)
@@
A.1.2. Calculating the order of convergence
@@
[…]
@@
A.2. Numerical examples
We now describe in more details the examples used to illustrate the order of convergence, except the non-homogeneous linear system, which is described in Section 8. We start with a linear equation driven by a Wiener process, then we illustrate the $\min\{H+1/2,1\}$ order of convergence in the case of a fractional Brownian motion (fBm) noise with Hurst parameter $0<H<1.$ Next we move to a number of linear and nonlinear models, with different semi-martingale noises, as mentioned in the Section 1.
A.2.1. Homogeneous linear equation with Wiener noise
@@
$\begin{cases}\displaystyle\frac{\mathrm{d}X_{t}}{\mathrm{d}t}=W_{t}X_{t},\qquad 0\leq t\leq T,\\ \left.X_{t}\right|_{t=0}=X_{0}.\end{cases}$ (A.2)
@@
Equation Equation A.2 has the explicit solution
$X_{t}=e^{\int_{0}^{t}W_{s}\;\mathrm{d}s}X_{0}.$ (A.3)
@@
The exact distribution of $\int_{0}^{\tau}W_{s}\;\mathrm{d}s$ given the step $\Delta W=W_{\tau}-W_{0}=W_{\tau}$ is computed in [20, Section 14.2] as
$\int_{0}^{\tau}W_{s}\;\mathrm{d}s\sim\frac{\tau}{2}\Delta W+\sqrt{\frac{\tau^{3}}{12}}\mathcal{N}(0,1)=\frac{\tau}{2}\Delta W+\mathcal{N}\left(0,\frac{\tau^{3}}{12}\right).$ (A.4)
As for the distribution of the integral over a mesh interval $t_{i},t_{i+1}]$ when given the endpoints $W_{t_{i}}$ and $W_{t_{i+1}},$ we use that $s\mapsto W_{t_{i}+s}-W_{t_{i}}$ is a standard Wiener process to find, from Equation A.4, that
@@
$Z_{i}\sim\mathcal{N}\left(0,\frac{(t_{i+1}-t_{i})^{3}}{12}\right)=\frac{\sqrt{(t_{i+1}-t_{i})^{3}}}{\sqrt{12}}\mathcal{N}(0,1).$ (A.5)
@@
$\int_{0}^{t_{j}}W_{s}\;\mathrm{d}s=\sum_{i=0}^{j-1}\int_{t_{i}}^{t_{i+1}}W_{s}\;\mathrm{d}s=\sum_{i=0}^{j-1}\left(\frac{(W_{t_{i+1}}+W_{t_{i}})}{2}(t_{i+1}-t_{i})+Z_{i}\right),$ (A.6)
where $Z_{0},\ldots,Z_{N-1}$ are independent random variables with distribution Equation A.5.
Notice the first term in Equation A.6 is the trapezoidal rule, with the second one being a correction term.
@@
Table 3. Mesh points (N), time steps (dt), strong error (error), and standard error (std err) of the Euler method for $\mathrm{d}X_{t}/\mathrm{d}t=W_{t}X_{t}$ for each mesh resolution NN, with initial condition $X_{0}\sim\mathcal{N}(0,1)$ and a standard Wiener process noise $\{W_{t}\}_{t}$, on the time interval $I=[0.0,1.0]$, based on $M=200$ sample paths for each fixed time step, with the target solution calculated with $65536$ points. The order of strong convergence is estimated to be $p=1.001$, with the 95% confidence interval $[0.936,1.0664]$.
Hence, once an Euler approximation of Equation A.2 is computed, along with realizations $\{W_{t_{i}}\}_{i=0}^{N}$ of a sample path of the noise, we consider an exact pathwise solution given by
$X_{t_{j}}=X_{0}e^{\sum_{i=0}^{j-1}\left(\frac{1}{2}\left(W_{t_{i}}+W_{t_{i+1}}\right)(t_{i+1}-t_{i})+Z_{i}\right)},$ (A.8)
for realizations $Z_{i}$ drawn from a normal distributions given by Equation A.5. Figure 10 shows an approximate solution and a few sample paths of exact solutions associated with the given realizations of the noise on the mesh points.
The function $f(t,x,y)=yx$ associated with equation Equation A.2 does not satisfy the uniform Lipschitz condition Equation 2.6. Nevertheless, thanks to the form of the exact solution Equation A.3 the form of the approximation Equation A.8 and estimates such as Equation A.7, the proof of Theorem 7.1 can be adapted to show strong order $1$ convergence in this case, as well.
@@
A.2.2. Fractional Brownian motion noise